# Do Spikes Protect Privacy? Investigating Black-Box Model Inversion Attacks in Spiking Neural Networks

### v1

@@ Abstract
As machine learning models become integral to security-sensitive applications, concerns over data leakage from adversarial attacks continue to rise. Model Inversion (MI) attacks pose a significant privacy threat by enabling adversaries to reconstruct training data from model outputs. While MI attacks on Artificial Neural Networks (ANNs) have been widely studied, Spiking Neural Networks (SNNs) remain largely unexplored in this context. Due to their event-driven and discrete computations, SNNs introduce fundamental differences in information processing that may offer inherent resistance to such attacks. A critical yet underexplored aspect of this threat lies in black-box settings, where attackers operate through queries without direct access to model parameters or gradients—representing a more realistic adversarial scenario in deployed systems. This work presents the first study of black-box MI attacks on SNNs. We adapt a generative adversarial MI framework to the spiking domain by incorporating rate-based encoding for input transformation and decoding mechanisms for output interpretation. Our results show that SNNs exhibit significantly greater resistance to MI attacks than ANNs, as demonstrated by degraded reconstructions, increased instability in attack convergence, and overall reduced attack effectiveness across multiple evaluation metrics. Further analysis suggests that the discrete and temporally distributed nature of SNN decision boundaries disrupts surrogate modeling, limiting the attacker’s ability to approximate the target model.
Index Terms:
Spiking Neural Networks, Black-Box Model Inversion Attacks, Neuromorphic Privacy, Adversarial Machine Learning
@@ I Introduction
Spiking Neural Networks (SNNs) have emerged as a promising computational paradigm, inspired by the biological processes of neuronal communication. By processing information through discrete spike events, SNNs offer a more biologically plausible and efficient alternative to traditional Artificial Neural Networks (ANNs) [1]. Their event-driven nature, in which neurons remain mostly inactive until a spike occurs, enables SNNs to excel in applications such as real-time object detection [2, 3], speech recognition[4], and robotics [5], where energy efficiency and temporal processing are crucial. However, as their use expands into critical areas, understanding and addressing their potential privacy risks is essential to maintain the integrity and confidentiality of sensitive information.
@@
Among the various privacy threats, Model Inversion (MI) attacks stand out due to their potential to reconstruct sensitive training data from a model’s outputs. These attacks exploit the learned patterns of the model to reverse-engineer sensitive information, often revealing training data or inferring private attributes [6]. MI attacks can be divided into two main categories: white-box and black-box. In white-box scenarios, attackers have full access to the model’s architecture and parameters, often employing gradient-based optimization for precise reconstructions. In contrast, black-box attacks operate under more restrictive conditions, relying solely on input-output queries to infer private information [7]. Given the real-world relevance of black-box scenarios, their impact on SNN privacy warrants deeper investigation [8].
@@
Recent advancements in black-box MI attacks have introduced diverse methodologies to extract sensitive information from machine learning models. A prominent approach involves generative models, such as Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs), which optimize in a latent space rather than directly in the input space [9, 10, 11]. These models iteratively adjust their latent variables to reconstruct candidate inputs that align with the target model’s outputs. Some methods further enhance the reconstruction quality by leveraging a learned prior from auxiliary public datasets for pre-training or fine-tuning [12, 13, 14]. While effective, this reliance on auxiliary datasets limits their applicability when such data are unavailable or do not align with the target domain. To address this, agnostic black-box methods have emerged, which operate independently of external datasets or assumptions about the target data distribution [15]. One such approach, GAMIN [16], combines adversarial optimization and surrogate modeling to infer sensitive inputs solely through direct queries to the target model.
@@
In SNNs, however, the prior studies have primarily explored privacy-preserving mechanisms and adversarial robustness. For example, differential privacy has been integrated into SNNs to protect training data by introducing noise during gradient updates, as demonstrated by [17]. Similarly, [18] addresses data and class leakage during ANN-to-SNN conversion through synthetic data generation and temporal learning rules. Other efforts, such as [19], have highlighted the relative resilience of SNNs against membership inference attacks due to their non-differentiable spiking dynamics. In the context of MI attacks, BrainLeaks [20] developed a tailored gradient-based attack to overcome the challenges of SNNs’ spiking mechanisms, demonstrating that while these models exhibit some resistance, they remain vulnerable to such privacy breaches. However, the privacy characteristics of SNNs under black-box MI attacks remain unexplored, leaving a significant gap in understanding their behavior in restricted-access scenarios.
@@ II-A Spiking Neural Networks (SNNs)
SNNs are biologically inspired computational models designed to emulate the temporal dynamics and communication mechanisms of biological neurons. Unlike ANNs, which rely on continuous-valued activations, SNNs process information via discrete, time-dependent spikes [21]. Each neuron in an SNN accumulates input spikes over time. When the membrane potential crosses a threshold, the neuron generates an output spike that propagates to downstream neurons. This event-driven communication makes SNNs energy-efficient and well-suited for neuromorphic applications.
Among various spiking neuron models, the Leaky Integrate-and-Fire (LIF) model is widely adopted due to its simplicity and biological relevance [22]. The membrane potential dynamics of a single LIF neuron in discrete time can be expressed as:
@@
For event-driven systems, input spikes can be directly obtained from neuromorphic sensors such as Dynamic Vision Sensors (DVS), which naturally generate spikes in response to changes in visual stimuli [23]. These sensors provide asynchronous spike-based data, making them inherently compatible with SNNs without requiring additional encoding.
However, to process non-spiking data in SNNs, inputs must be converted into spike trains using an encoding mechanism. One commonly used approach is rate encoding, where the spike frequency represents input intensity: higher input values correspond to higher spike rates, while lower values result in sparser spikes (see Figure 1). In this scheme, a continuous input value is mapped to a spike train by generating spikes with a probability proportional to the input magnitude over a fixed time window [24].
Training SNNs with backpropagation-based learning algorithms poses unique challenges due to the non-differentiable nature of spike-based activity. Particularly, the discontinuity of the spike generation function in Equation 2 blocks gradient flow, making standard backpropagation ineffective. To address this, one approach is to use surrogate gradient methods. These methods approximate the spiking function with a smooth, differentiable alternative during the backward pass, enabling techniques like Backpropagation Through Time (BPTT) to be applied effectively [25, 26].
@@ II-B Model Inversion (MI) Attacks
MI attacks aim to extract sensitive information from machine learning models by leveraging their learned representations. Formally, given a target model $T:\mathbb{R}^{d}\to\mathbb{R}^{k}$ and an observed output $\mathbf{y}=T(\mathbf{X})$ the objective of an MI attack is to approximate the input $\mathbf{X}\in\mathbb{R}^{d}$ that produced the output $\mathbf{y}\in\mathbb{R}^{k}$. This is often framed as an optimization problem [27]:
@@
where $\mathcal{L}(\cdot,\cdot)$ is a loss function measuring the similarity between the model’s output for a candidate input $\mathbf{X}$ and the target output $\mathbf{y}$. In white-box scenarios, attackers often leverage gradient information from the model to iteratively solve this optimization problem. Black-box scenarios, on the other hand, are more challenging, as they rely solely on input-output queries without access to internal parameters or gradients. One popular approach to tackle black-box MI attacks is to train a surrogate model that approximates the behavior of the target model. By querying the target model with synthetic inputs and observing its responses, the surrogate model learns to mimic the target model’s decision boundaries, enabling attackers to indirectly reconstruct input data[28].
@@ III Black-Box Model Inversion on SNNs
In this work, we utilize GAMIN (Generative Adversarial Model INversion) [16], a framework originally proposed for black-box model inversion on ANNs, and adapt it for the SNN domain. GAMIN works by training two neural networks in tandem: a generator $G$, which maps random noise $Z_{G}\sim\mathcal{N}(0,1)$ to synthetic inputs $X_{G}$, and a surrogate model $S$, which approximates the target model $T$. The goal is to generate inputs that, when passed through the surrogate, match the desired target label $y_{t}$ output by $T$. The training process consists of the following steps:
@@
To adapt GAMIN for SNNs, encoding and decoding mechanisms are utilized to handle the distinct input-output representations of these models. Rate encoding transforms static input data, such as pixel intensities, into spike trains by mapping intensities to spike rates over a fixed temporal window [29].
For output processing, decoding mechanisms convert spiking activity into confidence scores that GAMIN can process. Common decoding methods include spike count, time-averaged firing rate, and membrane potential [30].
@@ IV-A Experimental Setup
We evaluate the effectiveness of the attack on two tasks: face recognition and digit classification. For face recognition, we use the AT&T Face Database [31], which contains 400 grayscale images across 40 unique subjects. For digit classification, the MNIST [32] dataset is utilized, containing 70,000 grayscale images of handwritten digits. For SNNs, static image data from both datasets are converted into 25-step spiking representations using a rate-encoding scheme, where pixel intensities are mapped to spike rates.
The target models, both ANNs and SNNs, share the same fully connected architecture with a single hidden layer of 3,000 neurons to allow fair comparisons. For SNNs, we employ LIF neurons with a leakage rate of $0.7$. The SNNs are trained using backpropagation through time (BPTT) [25] with a fast-sigmoid function (slope of 40) for surrogate gradient calculations [26]. The loss function accumulates the softmax cross-entropy across all time steps, computed from the membrane potentials of the output neurons. Table I summarizes the test accuracy of the target models (ANNs and SNNs) trained on the MNIST and AT&T Face datasets.
We adopt the attack protocol from the GAMIN framework, which retains the best-performing model during training to prevent performance degradation from suboptimal updates, and reduce the impact of catastrophic forgetting. The query budget is set to 1,280,000 queries (or 20,000 batches of 64), aligning with the original framework to provide adequate query capacity for comprehensive evaluation. All implementations were performed using PyTorch [33] framework, with SNNtorch [30] employed for building and training the SNNs.
@@ IV-B Evaluation Metrics
We employ five metrics to analyze different aspects of the black-box MI attack and evaluate the performance of the generator, and surrogate models. Global Convergence Score ($M_{\text{global}}$), Fidelity Score ($F_{S}$), and Combined Accuracy ($A_{S\circ G}$) are adopted from GAMIN [16], while Surrogate Test Accuracy ($A_{S}$) and Target Model Accuracy on Inverted Samples ($A_{T}$) are introduced to provide additional insights into the surrogate and reconstructed samples.
@@ IV-B1 Global Convergence Score ($M_{\text{global}}$)
Inspired by the equilibrium loss in BEGAN [34], this metric assesses the overall performance of the attack by combining the surrogate model’s fidelity and the generator’s effectiveness. A lower $M_{\text{global}}$ value reflects improved alignment between the surrogate and generator and is defined as:
@@ V Discussion
Overall, these findings provide strong evidence that SNNs exhibit an inherent degree of robustness against black-box MI attacks. Unlike ANNs, where the generator can exploit smooth decision boundaries to refine reconstructions, the event-driven computations in SNNs appear to introduce irregularities that degrade the inversion process. However, the high surrogate fidelity on random inputs raises interesting questions about the nature of SNN decision boundaries and whether alternative encoding schemes (e.g., latency coding instead of rate coding) influence SNN privacy. Additionally, extending this analysis to neuromorphic datasets with real temporal dependencies, such as IBM DvsGesture [35], would provide a broader understanding of SNN privacy in real-world applications.
@@ VI Conclusion
Future research will investigate whether alternative spiking encodings, such as latency coding [29], impact inversion resistance, and assess attack performance on neuromorphic datasets with real temporal dependencies, such as event-based DVS datasets.

### v2

@@ Abstract
As machine learning models become integral to security-sensitive applications, concerns over data leakage from adversarial attacks continue to rise. Model Inversion (MI) attacks pose a significant privacy threat by enabling adversaries to reconstruct training data from model outputs. While MI attacks on Artificial Neural Networks (ANNs) have been widely studied, Spiking Neural Networks (SNNs) remain largely unexplored in this context. Due to their event-driven and discrete computations, SNNs introduce fundamental differences in information processing that may offer inherent resistance to such attacks. A critical yet underexplored aspect of this threat lies in black-box settings, where attackers operate through queries without direct access to model parameters or gradients, representing a more realistic adversarial scenario in deployed systems. This work presents the first study of black-box MI attacks on SNNs. We adapt a generative adversarial MI framework to the spiking domain by incorporating rate-based encoding for input transformation and decoding mechanisms for output interpretation. Our results show that SNNs exhibit significantly greater resistance to MI attacks than ANNs, as demonstrated by degraded reconstructions, increased instability in attack convergence, and overall reduced attack effectiveness across multiple evaluation metrics. Further analysis suggests that the discrete and temporally distributed nature of SNN decision boundaries disrupts surrogate modeling, limiting the attacker’s ability to approximate the target model.
@@ I Introduction
Spiking Neural Networks (SNNs) have emerged as a promising computational paradigm, inspired by the biological processes of neuronal communication. By processing information through discrete spike events, SNNs offer a more biologically plausible and efficient alternative to traditional Artificial Neural Networks (ANNs) [nunes2022spiking]. Their event-driven nature, in which neurons remain mostly inactive until a spike occurs, enables SNNs to excel in applications such as real-time object detection [cordone2022object, barchid2021deep], speech recognition[wu2020deep], and robotics [bing2018survey], where energy efficiency and temporal processing are crucial [poursiami2025scalable]. However, as their use expands into critical areas, understanding and addressing their potential privacy risks is essential to maintain the integrity and confidentiality of sensitive information[kim2022privatesnn, poursiami2024watermarking, tahmasivand2025lm, moshruba2025neuromorphic].
@@
Among the various privacy threats, Model Inversion (MI) attacks stand out due to their potential to reconstruct sensitive training data from a model’s outputs. These attacks exploit the learned patterns of the model to reverse-engineer sensitive information, often revealing training data or inferring private attributes [fredrikson2015model]. MI attacks can be divided into two main categories: white-box and black-box. In white-box scenarios, attackers have full access to the model’s architecture and parameters, often employing gradient-based optimization for precise reconstructions. In contrast, black-box attacks operate under more restrictive conditions, relying solely on input-output queries to infer private information [dibbo2023sok]. Given the real-world relevance of black-box scenarios, their impact on SNN privacy warrants deeper investigation [dionysiou2023exploring].
@@
Recent advancements in black-box MI attacks have introduced diverse methodologies to extract sensitive information from machine learning models. A prominent approach involves generative models, such as Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs), which optimize in a latent space rather than directly in the input space [nguyen2024label, han2023reinforcement, kahla2022label]. These models iteratively adjust their latent variables to reconstruct candidate inputs that align with the target model’s outputs. Some methods further enhance the reconstruction quality by leveraging a learned prior from auxiliary public datasets for pre-training or fine-tuning [khosravy2022model, xu2023sparse, ye2023c2fmi]. While effective, this reliance on auxiliary datasets limits their applicability when such data are unavailable or do not align with the target domain. To address this, agnostic black-box methods have emerged, which operate independently of external datasets or assumptions about the target data distribution [tramer2016stealing]. One such approach, GAMIN [aivodji2019gamin], combines adversarial optimization and surrogate modeling to infer sensitive inputs solely through direct queries to the target model.
@@
In SNNs, however, the prior studies have primarily explored privacy-preserving mechanisms and adversarial robustness. For example, differential privacy has been integrated into SNNs to protect training data by introducing noise during gradient updates, as demonstrated by [wang2022dpsnn]. Similarly, [kim2022privatesnn] addresses data and class leakage during ANN-to-SNN conversion through synthetic data generation and temporal learning rules. Other efforts, such as [moshruba2025neuromorphic, moshruba2025privacy, moshruba2025izhikevich], have highlighted the relative resilience of SNNs against membership inference attacks due to their non-differentiable spiking dynamics. In the context of MI attacks, BrainLeaks [poursiami2024brainleaks] developed a tailored gradient-based attack to overcome the challenges of SNNs’ spiking mechanisms, demonstrating that while these models exhibit some resistance, they remain vulnerable to such privacy breaches. However, the privacy characteristics of SNNs under black-box MI attacks remain unexplored, leaving a significant gap in understanding their behavior in restricted-access scenarios.
@@ II-A Spiking Neural Networks (SNNs)
SNNs are biologically inspired computational models designed to emulate the temporal dynamics and communication mechanisms of biological neurons. Unlike ANNs, which rely on continuous-valued activations, SNNs process information via discrete, time-dependent spikes [schuman2022opportunities]. Each neuron in an SNN accumulates input spikes over time. When the membrane potential crosses a threshold, the neuron generates an output spike that propagates to downstream neurons. This event-driven communication makes SNNs energy-efficient and well-suited for neuromorphic applications.
Among various spiking neuron models, the Leaky Integrate-and-Fire (LIF) model is widely adopted due to its simplicity and biological relevance [izhikevich2004model]. The membrane potential dynamics of a single LIF neuron in discrete time can be expressed as:
@@
For event-driven systems, input spikes can be directly obtained from neuromorphic sensors such as Dynamic Vision Sensors (DVS), which naturally generate spikes in response to changes in visual stimuli [hu2016dvs]. These sensors provide asynchronous spike-based data, making them inherently compatible with SNNs without requiring additional encoding.
However, to process non-spiking data in SNNs, inputs must be converted into spike trains using an encoding mechanism. One commonly used approach is rate encoding, where the spike frequency represents input intensity: higher input values correspond to higher spike rates, while lower values result in sparser spikes (see Figure 1). In this scheme, a continuous input value is mapped to a spike train by generating spikes with a probability proportional to the input magnitude over a fixed time window [gerstner2014neuronal].
Training SNNs with backpropagation-based learning algorithms poses unique challenges due to the non-differentiable nature of spike-based activity. Particularly, the discontinuity of the spike generation function in Equation 2 blocks gradient flow, making standard backpropagation ineffective. To address this, one approach is to use surrogate gradient methods. These methods approximate the spiking function with a smooth, differentiable alternative during the backward pass, enabling techniques like Backpropagation Through Time (BPTT) to be applied effectively [bellec2018long, neftci2019surrogate].
@@ II-B Model Inversion (MI) Attacks
MI attacks aim to extract sensitive information from machine learning models by leveraging their learned representations. Formally, given a target model $T:\mathbb{R}^{d}\to\mathbb{R}^{k}$ and an observed output $\mathbf{y}=T(\mathbf{X})$ the objective of an MI attack is to approximate the input $\mathbf{X}\in\mathbb{R}^{d}$ that produced the output $\mathbf{y}\in\mathbb{R}^{k}$. This is often framed as an optimization problem [struppek2022plug]:
@@
where $\mathcal{L}(\cdot,\cdot)$ is a loss function measuring the similarity between the model’s output for a candidate input $\mathbf{X}$ and the target output $\mathbf{y}$. In white-box scenarios, attackers often leverage gradient information from the model to iteratively solve this optimization problem. Black-box scenarios, on the other hand, are more challenging, as they rely solely on input-output queries without access to internal parameters or gradients. One popular approach to tackle black-box MI attacks is to train a surrogate model that approximates the behavior of the target model. By querying the target model with synthetic inputs and observing its responses, the surrogate model learns to mimic the target model’s decision boundaries, enabling attackers to indirectly reconstruct input data[fang2024privacy].
@@ III Black-Box Model Inversion on SNNs
In this work, we utilize GAMIN (Generative Adversarial Model INversion) [aivodji2019gamin], a framework originally proposed for black-box model inversion on ANNs, and adapt it for the SNN domain. GAMIN works by training two neural networks in tandem: a generator $G$, which maps random noise $Z_{G}\sim\mathcal{N}(0,1)$ to synthetic inputs $X_{G}$, and a surrogate model $S$, which approximates the target model $T$. The goal is to generate inputs that, when passed through the surrogate, match the desired target label $y_{t}$ output by $T$. The training process consists of the following steps:
@@
To adapt GAMIN for SNNs, encoding and decoding mechanisms are utilized to handle the distinct input-output representations of these models. Rate encoding transforms static input data, such as pixel intensities, into spike trains by mapping intensities to spike rates over a fixed temporal window [auge2021survey].
For output processing, decoding mechanisms convert spiking activity into confidence scores that GAMIN can process. Common decoding methods include spike count, time-averaged firing rate, and membrane potential [eshraghian2023training].
@@ IV-A Experimental Setup
We evaluate the effectiveness of the attack on two tasks: face recognition and digit classification. For face recognition, we use the AT&T Face Database [samaria1994orl], which contains 400 grayscale images across 40 unique subjects. For digit classification, the MNIST [deng2012mnist] dataset is utilized, containing 70,000 grayscale images of handwritten digits. For SNNs, static image data from both datasets are converted into 25-step spiking representations using a rate-encoding scheme, where pixel intensities are mapped to spike rates.
The target models, both ANNs and SNNs, share the same fully connected architecture with a single hidden layer of 3,000 neurons to allow fair comparisons. For SNNs, we employ LIF neurons with a leakage rate of $0.7$. The SNNs are trained using backpropagation through time (BPTT) [bellec2018long] with a fast-sigmoid function (slope of 40) for surrogate gradient calculations [neftci2019surrogate]. The loss function accumulates the softmax cross-entropy across all time steps, computed from the membrane potentials of the output neurons. Table I summarizes the test accuracy of the target models (ANNs and SNNs) trained on the MNIST and AT&T Face datasets.
We adopt the attack protocol from the GAMIN framework, which retains the best-performing model during training to prevent performance degradation from suboptimal updates, and reduce the impact of catastrophic forgetting. The query budget is set to 1,280,000 queries (or 20,000 batches of 64), aligning with the original framework to provide adequate query capacity for comprehensive evaluation. All implementations were performed using PyTorch [paszke2019pytorch] framework, with SNNtorch [eshraghian2023training] employed for building and training the SNNs.
@@ IV-B Evaluation Metrics
We employ five metrics to analyze different aspects of the black-box MI attack and evaluate the performance of the generator, and surrogate models. Global Convergence Score ($M_{\text{global}}$), Fidelity Score ($F_{S}$), and Combined Accuracy ($A_{S\circ G}$) are adopted from GAMIN [aivodji2019gamin], while Surrogate Test Accuracy ($A_{S}$) and Target Model Accuracy on Inverted Samples ($A_{T}$) are introduced to provide additional insights into the surrogate and reconstructed samples.
@@ IV-B1 Global Convergence Score ($M_{\text{global}}$)
Inspired by the equilibrium loss in BEGAN [berthelot2017began], this metric assesses the overall performance of the attack by combining the surrogate model’s fidelity and the generator’s effectiveness. A lower $M_{\text{global}}$ value reflects improved alignment between the surrogate and generator and is defined as:
@@ V Discussion
Overall, these findings provide strong evidence that SNNs exhibit an inherent degree of robustness against black-box MI attacks. Unlike ANNs, where the generator can exploit smooth decision boundaries to refine reconstructions, the event-driven computations in SNNs appear to introduce irregularities that degrade the inversion process. However, the high surrogate fidelity on random inputs raises interesting questions about the nature of SNN decision boundaries and whether alternative encoding schemes (e.g., latency coding instead of rate coding) influence SNN privacy. Additionally, extending this analysis to neuromorphic datasets with real temporal dependencies, such as IBM DvsGesture [amir2017low], would provide a broader understanding of SNN privacy in real-world applications.
@@ VI Conclusion
Future research will investigate whether alternative spiking encodings, such as latency coding [auge2021survey], impact inversion resistance, and assess attack performance on neuromorphic datasets with real temporal dependencies, such as event-based DVS datasets.
@@
Acknowledgment
The research was funded in part by National Science Foundation through award CCF2319619 and Department of Energy through award DE-SC0025349.